OpenStudy (anonymous):

Help? & Explain please :( A community garden group is giving boxes of plants to local gardeners. The people who assembled the boxes randomly picked items from 3 kinds of vegetable plants, 3 kinds of trees, 2 types of flowers, and 5 types of herbs. Find the probability that Greg will pick a box with a cauliflower plant, a white ash tree, a daylily, and a tarragon herb plant.

OpenStudy (anonymous):

\[\frac{1}{3}\times \frac{1}{3}\times \frac{1}{2}\times \frac{1}{5}\]

OpenStudy (anonymous):

That equals 1/90

OpenStudy (anonymous):

I'm confused :(

OpenStudy (amistre64):

the probability of picking a "group" is what .. 3/13 , 3/13 ,2/13, and 5/13 but when we are picking a specific item out of the groups; we simply multiply the specific probabilites of each group as satellite has given

OpenStudy (amistre64):

i believe it has to do with a condition of "given that" but I cant seem to parse it out that well

OpenStudy (anonymous):

yes \[\frac{1}{90}\] is the answer

OpenStudy (amistre64):

1/13 1/13 1/13 1/13 ---- * ----- * ----- * ----- ; are the probabilities, but the (/13)s 3/13 3/13 2/13 5/13 cancel each other out